OpenStudy (anonymous):

Find p'(2) If p(x)=3^2x

OpenStudy (nirmalnema):

replace x by 2 in the que.... so u will get 3^4 means 3*3*3*3 =81

OpenStudy (anonymous):

Nooooooo. its wrong. there is a prime there!

OpenStudy (nirmalnema):

prime there means

OpenStudy (loser66):

derivative!! @nirmalnema

OpenStudy (anonymous):

p'(x) = 2*3^(2x).ln3 p'(2) = 2*3^(2*2).ln3=2*81*ln3=162.ln3

OpenStudy (nirmalnema):

ohh..okk

OpenStudy (anonymous):

Or p(x) = 9^x p'(x) = 9^x*ln9 = 2*9^x*ln3 p'(2) = 2*9^2*ln3= 162*ln3 again if it is hard to derivate something^2x

OpenStudy (anonymous):

Thanks so much @zarax

OpenStudy (anonymous):

You're welcome